OpenStudy (anonymous):

It has been estimated that the number of Venutian Carp in the Tethis River is given by V(t)= 3t^2+8t+203 Where t is the number of years since 1990 A.Evaluate V'(5) B. Interpret what V'(5)means

OpenStudy (anonymous):

What's stopping you?

OpenStudy (anonymous):

what do you mean whats stopping me

OpenStudy (anonymous):

Can't you do it on your own?

OpenStudy (anonymous):

I am confussed on part A on where to begin

OpenStudy (anonymous):

Do you know how to find \(V'(t)\)?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Well\[ V'(t) = \lim_{h\to 0}\frac{V(t+h)-V(t)}{h}\\ \]Which means: \[\begin{split} V'(5) &= \lim_{h\to 0}\frac{V(5+h)-V(5)}{h}\\ &=\lim_{h\to 0}\frac{3(5+h)^2+8(5+h)+203-[3(5)^2+8(5)+203]}{h} \end{split}\] Good luck =)

OpenStudy (anonymous):

It would be really nice if you knew how to differentiate, wouldn't it?

OpenStudy (anonymous):

yes

OpenStudy (dan815):

LOL wio giving first principle ><

OpenStudy (anonymous):

Try to simplify that fraction a bit @farmergirl411

OpenStudy (anonymous):

ok give me a minute

OpenStudy (anonymous):

ok I got ((9h^2+168h+636/h))

OpenStudy (anonymous):

is that right

OpenStudy (anonymous):

No it is not.

OpenStudy (anonymous):

what didi I do wrong

OpenStudy (anonymous):

\[ \begin{split} V'(5) &= \lim_{h\to 0}\frac{V(5+h)-V(5)}{h}\\ &=\lim_{h\to 0}\frac{3(5+h)^2+8(5+h)+203-[3(5)^2+8(5)+203]}{h}\\ &=\lim_{h\to 0}\frac{3(5+h)^2+8(5+h)-[3(5)^2+8(5)]}{h}\\ &=\lim_{h\to 0}\frac{3(5)^2+3(10h)+3(h)^2+8(5)+8h-[3(5)^2+8(5)]}{h}\\ &=\lim_{h\to 0}\frac{3(10h)+3(h)^2+8h}{h}\\ &=\lim_{h\to 0}\frac{30h+3h^2+8h}{h}\\ &=\lim_{h\to 0}30+3h+8\\ &=30+8\\ &=38 \end{split} \]

OpenStudy (anonymous):

I don't know where you went wrong, but wherever you didn't match me, you must have gone down the wrong path.

OpenStudy (anonymous):

must be so then 38 is the anwer for part a

OpenStudy (dan815):

203−[3(5)2+8(5)+203] = -75-40 = -115

OpenStudy (anonymous):

I am confussed who is right

OpenStudy (dan815):

im just showing u that u messed up the calculations for the constants

OpenStudy (anonymous):

oh ok but so is the 38 right that wio got

OpenStudy (dan815):

6t+8= v'(5) = 38

OpenStudy (dan815):

yah

OpenStudy (anonymous):

ok so you by any chance know how to do part b of this problem

OpenStudy (dan815):

you should practice simplying the first principle tho it usually cancels when you expand all the brackets!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do you know how to do part b

OpenStudy (mertsj):

Do you know that the first derivative gives you the slope or rate of change? In this case, the rate of change interpretation is more helpful. It tells you the rate of change in the number of carp